1L flask contains nitrogen along with a drop or two of water at 40° C, the total pressure being 760 torr. If all the contents are transferred to another flask of 0.5L at the same temperature, the pressure set up in the second flask will be (aqueous tension at 40°C = 55 torr)


 



(a)   1410 torr


(b)   1465 torr


(c)   1520 torr


(d)   None of these

Grade:12

2 Answers

GOVINDARAJ IYYAPPAN
13 years ago

AS PER BOYLE'S  LAW 'PRESSURE INVERSELY VARIES WITH VOLUME' FOR A GIVEN MASS OF GAS AT FIXED TEMPERATURE.

PRESSURE OF DRY NITROGEN IS 760-55=705 TORR.SINCE THE VOLUME IS REDUCED TO ONE HALF THE PRESSURE WILL BE DOUBLED. WHICH IS 705X2=1410 TORR.

OR IT CAN BE SOLVED USING THE FORMULA

P1V1=P2V2

705X1=P2X0.5

P2=705/0.5

  =1410 torr

sn sharma
12 years ago

using Boyle's law,final pressure exerted by nitrogen alone can be calculated as 1410 torr.

And then by adding saturated water vapour pressure at 40degree celcius,we get..

  1410+55 =1465 torr.

Hence the correct answer for this is B.